OpenStudy (help_people):

The following graph describes function 1, and the equation below it describes function 2: Function 1 graph of function f of x equals negative x squared plus 8 multiplied by x minus 15 Function 2 f(x) = −x2 + 2x − 3 Function ____ has the larger maximum. (Put 1 or 2 in the blank space) Numerical Answers Expected! Answer for Blank 1:

OpenStudy (lfreeman1285):

Your first step is to write out equation 1. Then you have to find the vertex of each equation, x = \[\frac{ -b }{ 2a }\] Then plug that value back into its respective equation. Whichever one produces the highest number is the one with the greater maximum.

OpenStudy (lfreeman1285):

So for the first equation: f(x)=-x2+8x-15 -b/2a = -8/-2 = 4 -(4)2 + 8(4) -15 = 1 For the second equation: -b/2a = -2 / -2 = 1 -(1)2 +2(1) - 3 = -2 1 > -2 And therefore, equation 1 has the higher maximum

OpenStudy (help_people):

ty!